# CurvFed: Curvature-Aligned Federated Learning for Fairness without Demographics

### v6

@@ 3.1. Problem Context and Literature Gap
Each client $t$ trains a local model using $\mathcal{D}_{t}$ with parameters $\theta^{*}_{t}$ by minimizing an empirical risk function over its dataset:
@@
After local training, the server aggregates the client models $\{\theta_{1}^{*},\dots,\theta_{q}^{*}\}$ into a global model with parameter $\theta^{*}$. This aggregation is typically done through a weighted average, as shown in Eq. 4, where each client is assigned a non-negative weight $\beta_{t}$ such that $\sum_{t=1}^{q}\beta_{t}=1$, with the goal of minimizing its excessive loss (papadaki2024federated, ), shown in Eq. 5.
@@
A key fairness concern is that the global model may disproportionately favor some clients. To address this, prior work (papadaki2024federated, ) introduces the concept of excessive loss in Eq. 5—the performance gap between a client’s local and the global model—and promotes fairness by minimizing its variance across clients:
@@
This objective ensures that model utility is distributed equitably across clients. In this work, we utilize $\|.\|$ notation to denote the Euclidean norm, vectors or matrices by bold symbols, and sets by uppercase symbols.
@@ 3.2. FWD in FL: A Curvature-Based Roadmap
(7) $R(t)\leq\|g_{t}^{\ell}\|\times\|\theta^{*}-\theta^{*}_{t}\|+\frac{1}{2}\lambda(\textbf{H}_{t}^{\ell})\times\|\theta^{*}-\theta^{*}_{t}\|^{2}+\mathcal{O}(\|\theta^{*}-\theta^{*}_{t}\|^{3}).$
Here, $g_{t}^{\ell}=\nabla_{\theta}\ell(\theta^{*}_{t},\mathcal{D}_{t})$ is the local models gradient vector, and $\textbf{H}_{t}^{\ell}=\nabla^{2}_{\theta}\ell(\theta^{*}_{t},\mathcal{D}_{t})$ is the local model’s Hessian at optimal parameter $\theta^{*}_{t}$ associated with the loss function $\ell$. The term $\lambda(\cdot)$ denotes the maximum eigenvalue (hereby referenced as the top eigenvalue). Equation 7 reveals that the excessive loss is tightly linked to the client $t$’s (i) gradient vector corresponding to the loss, i.e., $g_{t}^{\ell}$ and (ii) the top Hessian eigenvalue, a proxy of the curvature (sharpness) of the local loss landscape.
Sharpness, Curvature, and Fairness: High curvature in the loss landscape of model optimization, characterized by sharp minima, typically associated with large top Hessian eigenvalues—is known to impair generalization (dauphin2024neglected, ; jastrzkebski2018relation, ). Importantly, disparities in two groups (based on demography or other factors) in terms of their top Hessian eigenvalues have been shown to correlate with the model’s performance disparities between those groups (tran2022pruning, ), suggesting that mitigating disparities among all groups’ curvatures (top Hessian eigenvalues) will promote fairness.
@@
    Computing the full Hessian matrix is computationally costly (dong2017learning, ), thus, infeasible on resource-constrained edge devices.
@@
Lee et. al. (lee2022masking, ) describe the relationship between the empirical FIM (F) and the Hessian matrix (H) as shown in equation 8. Where $\nabla_{\theta}\ell(f_{\theta}(x_{i}),y_{i})$ is the the gradient vector corresponding to the loss of the $i^{th}$ data sample, for a model $f$ with parameters $\theta$ and the total number of data samples represented by $n_{t}$. The empirical FIM thus serves as a computationally efficient approximation of the Hessian, reducing computational costs during training (singh2020woodfisher, ; lee2022masking, ).
@@
Using Jensen’s and Weyl’s inequalities (boyd2004convex, ; horn2012matrix, ), with an assumption of approximate dominant eigenvector alignment for the lower bound, we establish bounds:
@@
Here, $g_{t}^{\ell}$ is the local model’s gradient at its optimum $\theta_{t}^{*}$, and $\lambda(\mathbf{F}_{t}^{\ell})$ is the top eigenvalue of its FIM, serving as a proxy for Hessian’s top eigenvalue (following Proposition 1). This indicates that for a client $t$, its excessive loss is impacted mostly by $g_{t}^{\ell}$ and $\lambda(\mathbf{F}_{t}^{\ell})$, which may vary significantly across clients. Our variance reduction strategy builds on this observation. During its local training, each client $t$ seeks to minimize its empirical risk $J(\theta,\mathcal{D}_{t})$. Successful training results in a small local gradient norm $\|g_{t}^{\ell}\|$ at $\theta_{t}^{*}$. Furthermore, as established in Proposition 2, our framework encourages clients to find solutions $\theta_{t}^{*}$ that also exhibit low curvature, i.e., a minimized $\lambda(\mathbf{F}_{t}^{\ell})$. That means an effective locally trained client will have lower $g_{t}^{\ell}$ and $\lambda(\mathbf{F}_{t}^{\ell})$, thus lower excessive loss, and the opposite holds for ineffective locally trained ones.
@@ 3.3. Assumptions
    We assume the central server is not resource-constrained and, consistent with standard FL approaches (e.g., FedAvg(mcmahan2017communication, )), can perform model aggregation efficiently.
@@ 4. CurvFed: Design Choices and Approach
Figure 1. CurvFed overview.
@@ 4.1.1. Fair Local Training Objective:
In equation 11, we introduce $\alpha$, a hyperparameter that acts as a balancing factor between the two objectives. $\alpha$ is identified using a grid search in this work. However, popular hyperparameter optimization frameworks like Optuna (akiba2019optuna, ) can be used to find the optimal $\alpha$. In our evaluations, a single $\alpha$ identified via grid search was effective across all three datasets. This consistency suggests that the local optimization strategy transfers reasonably well across heterogeneous human-sensing datasets. We will use $\lambda(F_{client})$ instead of $\lambda(F^{\ell_{\text{client}}})_{\text{correct}}$ for simplicity for the rest of the section.
@@ 4.1.2. Sharpness-aware Aggregation to Promote Fairness:
This aggregation strategy draws inspiration from Pessimistic Weighted Aggregation (P-W) in FL (djebrouni2022towards, ), which excludes clients with known biases during aggregation. However, in the FWD setting, such bias information is unavailable. Instead, our Sharpness-aware Aggregation implicitly promotes fairness by assigning higher weights to clients with lower excessive loss, i.e., having lower top eigenvalue of FIM and lower classification error (according to Proposition 3), thus supporting fairer and more performant global updates without relying on explicit sensitive attributes.
@@ 4.2. Approach
Algorithm 1 outlines CurvFed approach, which begins with the following inputs: a randomly initialized global model denoted as $G^{0}$, the number of clients participating in FL process represented by N, the total number of rounds for the FL process indicated by R, the number of epochs E for conducting local training on each client’s side, the cycle length c for implementing the stochastic weight averaging (SWA) protocol (caldarola2022improving, ), the weighting factor $\alpha$ used to balance the client loss, the batch size B utilized by the clients during training, client learning rate $\gamma$ and learning rate scheduler $\gamma()$. Different modules of CurvFed are detailed below:
@@ 4.2.1. Local Client Training:
Following standard FL protocol (mcmahan2017communication, ), the server initializes by distributing a randomly initialized global model $G^{0}$ and other training parameters to the clients for local training through the $TrainClient$ module (line 12-13).
TrainClient Module, as shown in Algorithm 2, receives the global model, along with training configurations such as the number of local training epochs $E$, batch size $B$, learning rate $\gamma$, and the client loss weighting factor $\alpha$ from the server. On the client side, data is divided into a training set $\mathcal{D}_{train}$ and an evaluation set $\mathcal{D}_{eval}$. The model is trained on $\mathcal{D}_{train}$ (line-9), where we utilized the Sharpness-Aware Minimization (SAM) optimizer for model training(foret2021sharpnessaware, ). The choice of the model optimizer is influenced by prior works in FL, which claim that using SAM optimizer at the client level helps with convergence and promotes local flatness (caldarola2022improving, ), thus lowering curvature and further promoting fairness.
@@ 4.2.2. Sharpness-aware Aggregation:
Stochastic Weight Averaging (SWA): On the server side, we also use SWA (izmailov2018averaging, ), a technique shown to enhance generalization in FL systems (caldarola2022improving, ). The significance of generalization in fair models is underscored by research (ferry2023improving, ; cotter2018training, ) focusing on improving their robustness and ability to perform fairly across unseen data. This addresses the challenge of fairness constraint overfitting (cotter2018training, ), where models fair on training data may still exhibit unfairness on unseen data. In human-sensing FL, ensuring fairness and robustness across unseen instances is critical; hence, the adoption of SWA is crucial in CurvFed. SWA averages the global models at regular intervals, i.e., cycles $c$ (line 19), and adjusts the learning rate throughout the FL process (line 9) to achieve flatter minima, thereby preventing overfitting as the process advances. The SWA process begins after a threshold ($\eta$) number of rounds are complete (lines 4-7); for our evaluations, this threshold is set to $\eta=20\%$. The early start threshold was chosen due to the limited data size, which allows for quicker convergence in our setting.
@@ 5. Experimental Evaluation
This section provides a comprehensive evaluation of CurvFed, covering datasets and models (Section 5.1), followed by an analysis addressing four key research questions on CurvFed’s effectiveness (Section 5.2). We then assess its real-world deployment feasibility and sensitivity (Sections 5.3 and 5.4) and conclude with an ablation study justifying the design choices (Section 5.5).
5.1. Datasets and Models
Few human-sensing datasets capture diverse bias-inducing factors such as sex and sensor placement (yfantidou2023beyond, ). We evaluate CurvFed on three such datasets: Stress Sensing (xiao2024reading, ) (48 participants), WIDAR (zhang2021WIDAR3, ) (16 participants), and HugaDB (chereshnev2018hugadb, ) (18 participants). Each includes known sources of bias—WIDAR: sex and sensor orientation; Stress Sensing: sex and watch hand; HugaDB: sex.
Dataset Rationale: We selected these datasets to evaluate the effectiveness of CurvFed across diverse human-sensing applications:
Human Activity Recognition (HAR): WIDAR and HugaDB utilize different sensing modalities— WiFi-based body-coordinate velocity profiles (zhang2021WIDAR3, ) in WIDAR and inertial data (accelerometer and gyroscope) in HugaDB—for gesture and activity recognition, respectively.
Stress Sensing: The Stress sensing dataset includes multi-modal input data, specifically electrodermal activity (EDA) and accelerometer readings from an Empatica E4 watch for stress detection.
This selection ensures a comprehensive evaluation of CurvFed across different modalities and application domains.
Model Architectures: We use task-specific MLPs: WIDAR (2-layer, 40% dropout), HugaDB (3-layer, 30% dropout with instance norm), and Stress Sensing (3-layer, batch norm). We apply an 80:20 train-test split per client and adapt benchmark models (yang2023benchmark, ; sharma2022psychophysiological, ; pirttikangas2006feature, ) for FL classification.

### v7

@@ 3.1. Problem Context and Literature Gap
Each client $t$ trains a local model using $\mathcal{D}_{t}$ with parameters $\theta^{*}_{t}$ by minimizing an empirical risk function $J(\theta,\mathcal{D}_{t})$, which represents the average empirical loss over the local dataset,
@@
After local training, the server aggregates the client models $\{\theta_{1}^{*},\dots,\theta_{q}^{*}\}$ into a global model with parameter $\theta^{*}$. This aggregation is typically done through a weighted average, as shown in Eq. 4, where each client is assigned a non-negative weight $\beta_{t}$ such that $\sum_{t=1}^{q}\beta_{t}=1$, with the goal of minimizing its excessive loss ($R$) (Papadaki et al., 2024), shown in Eq. 5.
@@
Here, $\theta_{t}^{*}$ denotes the locally optimal model parameters obtained by client $t$, whereas $\theta^{*}$ represents the aggregated global model parameters produced by the central server. The excessive loss $R(t)$ quantifies the performance degradation experienced by client $t$ when using the global model instead of its locally optimized model, evaluated on $\mathcal{D}_{t}$.
A key fairness concern is that the global model may disproportionately favor some clients. To address this, prior work (Papadaki et al., 2024) introduces the concept of excessive loss in Eq. 5—the performance gap between a client’s local and the global model—and promotes fairness by minimizing its variance across clients :
@@
where, the operator $\text{Var}_{t}[\cdot]$ denotes variance computed across all clients indexed by $t\in\{1,\dots,q\}$.
This objective ensures that model utility is distributed equitably across clients. Throughout this work, $\|\cdot\|$ denotes the Euclidean norm, vectors and matrices are represented by bold symbols, and sets are denoted by uppercase letters, unless stated otherwise.
@@ 3.2. FWD in FL: A Curvature-Based Roadmap
(7) $R(t)\leq\|g_{t}^{\ell}\|\times\|\theta^{*}-\theta^{*}_{t}\|+\frac{1}{2}\lambda(\textbf{H}_{t}^{\ell})\times\|\theta^{*}-\theta^{*}_{t}\|^{2}+O(\|\theta^{*}-\theta^{*}_{t}\|^{3}).$
Here, $g_{t}^{\ell}=\nabla_{\theta}\ell(\theta^{*}_{t},\mathcal{D}_{t})$ is the local models gradient vector, and $\textbf{H}_{t}^{\ell}=\nabla^{2}_{\theta}\ell(\theta^{*}_{t},\mathcal{D}_{t})$ is the local model’s Hessian at optimal parameter $\theta^{*}_{t}$ associated with the loss function $\ell$. The term $\lambda(\cdot)$ denotes the maximum eigenvalue (hereby referenced as the top eigenvalue) and the big-oh term $O(\|\theta^{*}-\theta^{*}_{t}\|^{3})$ represents all the higher order terms in the expansion. Equation 7 reveals that the excessive loss is tightly linked to the client $t$’s (i) gradient vector corresponding to the loss, i.e., $g_{t}^{\ell}$ and (ii) the top Hessian eigenvalue, a proxy of the curvature (sharpness) of the local loss landscape.
Sharpness, Curvature, and Fairness: High curvature in the loss landscape of model optimization, characterized by sharp minima, typically associated with large top Hessian eigenvalues—is known to impair generalization (Dauphin et al., 2024; Jastrzębski et al., 2018). Importantly, disparities in two groups (based on demography or other factors) in terms of their top Hessian eigenvalues have been shown to correlate with the model’s performance disparities between those groups (Tran et al., 2022), suggesting that mitigating disparities among all groups’ curvatures (top Hessian eigenvalues) will promote fairness.
@@
    Computing the full Hessian matrix is computationally costly (Dong et al., 2017), thus, infeasible on resource-constrained edge devices.
@@
For notational simplicity, $\mathbf{H}$ here denotes the client-specific Hessian $\mathbf{H}_{t}^{\ell}$ evaluated at $\theta_{t}^{*}$.
Lee et. al. (Lee et al., 2022) describe the relationship between the empirical FIM (F) and the Hessian matrix (H) as shown in equation 8. Where $\nabla_{\theta}\ell(f_{\theta}(x_{i}),y_{i})$ is the the gradient vector corresponding to the loss of the $i^{th}$ data sample, for a model $f$ with parameters $\theta$ and the total number of data samples represented by $n_{t}$. The empirical FIM thus serves as a computationally efficient approximation of the Hessian, reducing computational costs during training (Singh and Alistarh, 2020; Lee et al., 2022).
@@
Using Jensen’s and Weyl’s inequalities (Boyd and Vandenberghe, 2004; Horn and Johnson, 2012), with an assumption of approximate dominant eigenvector alignment for the lower bound, we establish bounds:
@@
Here, The quantity $\delta_{t}$ measures the parameter-space deviation between the global model and client $t$’s locally optimal model. $g_{t}^{\ell}$ is the local model’s gradient at its optimum $\theta_{t}^{*}$, and $\lambda(\mathbf{F}_{t}^{\ell})$ is the top eigenvalue of its FIM, serving as a proxy for Hessian’s top eigenvalue (following Proposition 1). This indicates that for a client $t$, its excessive loss is impacted mostly by $g_{t}^{\ell}$ and $\lambda(\mathbf{F}_{t}^{\ell})$, which may vary significantly across clients. Our variance reduction strategy builds on this observation. During its local training, each client $t$ seeks to minimize its empirical risk $J(\theta,\mathcal{D}_{t})$. Successful training results in a small local gradient norm $\|g_{t}^{\ell}\|$ at $\theta_{t}^{*}$. Furthermore, as established in Proposition 2, our framework encourages clients to find solutions $\theta_{t}^{*}$ that also exhibit low curvature, i.e., a minimized $\lambda(\mathbf{F}_{t}^{\ell})$. That means an effective locally trained client will have lower $g_{t}^{\ell}$ and $\lambda(\mathbf{F}_{t}^{\ell})$, thus lower excessive loss, and the opposite holds for ineffective locally trained ones.
@@
3.3. Synthesis of Theoretical Insights: Curvature Alignment as a Proxy for Fairness Without Demographics (FWD) in Human-centric federated learning (FL)
The above theoretical results motivate curvature alignment as a principled and operational mechanism for achieving fairness under the challenging setting of FWD in human-centric FL.
As discussed in Section 2.1, conventional fairness-attaining approaches rely on explicit group-based metrics (e.g., the Equal Opportunity gap), which require access to sensitive or bias-inducing attribute labels. In our FWD setting, such labels are unavailable during training, and FL privacy constraints prohibit their use. Therefore, standard group-equalization strategies are not implementable. Moreover, human-sensing systems often involve multiple simultaneous bias-inducing factors (e.g., sex, body type, sensor placement) (Jui and Rivas, 2024), and mitigating one explicitly may exacerbate another (Djebrouni et al., 2024). Therefore, this paper’s challenge becomes more complex: ensuring fairness under unknown and potentially multi-attribute biases, rather than across predefined groups with known sensitive attributes.
Under these constraints, we pursue an alternative pathway grounded in recent ML optimization theories linking loss-landscape geometry to group-wise performance disparities. As Section 3, establishes that a client’s local model’s loss landscape curvature being flatter reduces disparities in generalization performance across groups within that client’s local data (within client disparity reduction established through Theorem 1 and Proposition 1), and the loss landscape curvatures being similar across clients’ local models reduce disparities in generalization performance across clients’ local data (inter-client established in Proposition 2). Thus, loss landscape curvature is not treated as fairness itself, but as a measurable, optimization-level quantity that upper bounds performance disparity in the absence of information about demographic or bias-creating factors.
In summary, by promoting (1) Flatter curvature within clients, we reduce intra-client (even unknown) group-wise disparity, and (2) Similar curvature across clients, we reduce inter-client local data performance disparities. These together provide an attribute-agnostic pathway toward fairness across groups, whether a single or multiple sensitive or demographic factors are creating the disparity.
Importantly, curvature information (via the FIM top eigenvalue) does not reveal demographic information and can be computed locally, satisfying the constraints of privacy-preserving fairness without demographic (FWD) scope of FL. This makes curvature geometric consistency a feasible operational mechanism to attain FWD, whereas conventional fairness mechanisms are not implementable under the same constraints.
@@ 3.4. Assumptions
    We assume the central server is not resource-constrained and, consistent with standard FL approaches (e.g., FedAvg(McMahan et al., 2017)), can perform model aggregation efficiently.
  4. (4)
    We assume that clients permit sharing loss-landscape curvature statistics (e.g., the top FIM eigenvalue) during FL aggregation. These statistics capture aggregate optimization geometry and do not involve sharing raw data, gradients, or sensitive attributes. To the best of our knowledge, there is no evidence that such curvature summaries can be used to reconstruct demographic or sensitive attribute information; thus, satisfying the FWD setting.
@@ 4. CurvFed: Design Choices and Approach
Figure 1. Overview of CurvFed, illustrating client-side training, sharpness-aware aggregation, and global model updates in federated learning.
@@ 4.1.1. Fair Local Training Objective:
In equation 11, we introduce $\alpha$, a hyperparameter that acts as a balancing factor between the two objectives. $\alpha$ is identified using a grid search in this work. However, popular hyperparameter optimization frameworks like Optuna (Akiba et al., 2019) can be used to find the optimal $\alpha$. In our evaluations, the $\alpha$ values identified via grid search were effective across all three datasets, with one dataset exhibiting an optimal value within a similar range. A detailed discussion of the sensitivity to the weighting factor is provided in Section 6.2.4. This consistency suggests that the local optimization strategy transfers reasonably well across heterogeneous human-sensing datasets. We will use $\lambda(F_{client})$ instead of $\lambda(F^{\ell_{\text{client}}})_{\text{correct}}$ for simplicity for the rest of the section.
@@ 4.1.2. Sharpness-aware Aggregation to Promote Fairness:
This aggregation strategy draws inspiration from Pessimistic Weighted Aggregation (P-W) in FL (Djebrouni, 2022), which excludes clients with known biases during aggregation. However, in the FWD setting, such bias information is unavailable. Instead, our Sharpness-aware Aggregation implicitly promotes fairness by assigning higher weights to clients with lower excessive loss, i.e., having lower top eigenvalue of FIM and lower classification error (according to Proposition 3), thus supporting fairer and more performant global updates without relying on explicit sensitive attributes.
@@ 4.2. Approach
Algorithm 1 outlines CurvFed approach, which begins with the following inputs: a randomly initialized global model denoted as $G^{0}$, the number of clients participating in FL process represented by N, the total number of rounds for the FL process indicated by R, the number of epochs E for conducting local training on each client’s side, the cycle length c for implementing the stochastic weight averaging (SWA) protocol (Caldarola et al., 2022), the weighting factor $\alpha$ used to balance the client loss, the batch size B utilized by the clients during training, client learning rate $\gamma$ and learning rate scheduler $\gamma()$. Different modules of CurvFed are detailed below:
@@ 4.2.1. Local Client Training:
Following standard FL protocol (McMahan et al., 2017), the server initializes by distributing a randomly initialized global model $G^{0}$ and other training parameters to the clients for local training through the $TrainClient$ module (line 12-13).
TrainClient Module, as shown in Algorithm 2, receives the global model, along with training configurations such as the number of local training epochs $E$, batch size $B$, learning rate $\gamma$, and the client loss weighting factor $\alpha$ from the server. On the client side, data is divided into a training set $\mathcal{D}_{train}$ and an evaluation set $\mathcal{D}_{eval}$. The model is trained on $\mathcal{D}_{train}$ (line-9), where we utilized the Sharpness-Aware Minimization (SAM) optimizer for model training(Foret et al., 2021). The choice of the model optimizer is influenced by prior works in FL, which claim that using SAM optimizer at the client level helps with convergence and promotes local flatness (Caldarola et al., 2022), thus lowering curvature and further promoting fairness.
@@ 4.2.2. Sharpness-aware Aggregation:
Stochastic Weight Averaging (SWA): On the server side, we also use SWA (Izmailov et al., 2018), a technique shown to enhance generalization in FL systems (Caldarola et al., 2022). The significance of generalization in fair models is underscored by research (Ferry et al., 2023; Cotter et al., 2018) focusing on improving their robustness and ability to perform fairly across unseen data. This addresses the challenge of fairness constraint overfitting (Cotter et al., 2018), where models fair on training data may still exhibit unfairness on unseen data. In human-sensing FL, ensuring fairness and robustness across unseen instances is critical; hence, the adoption of SWA is crucial in CurvFed. SWA averages the global models at regular intervals, i.e., cycles $c$ (line 19), and adjusts the learning rate throughout the FL process (line 9) to achieve flatter minima, thereby preventing overfitting as the process advances. The SWA process begins after a threshold ($\eta$) number of rounds are complete (lines 4-7); for our evaluations, this threshold is set to $\eta=20\%$. The early start threshold was chosen due to the limited data size, which allows for quicker convergence in our setting.
@@ 5. Experimental Evaluation
This section provides a comprehensive evaluation of CurvFed, covering datasets and models (Section 5.1), followed by an analysis addressing four key research questions on CurvFed’s effectiveness (Section 5.3). In this work, we compared CurvFed against standard benign FL models—FedAvg(McMahan et al., 2017), FedSAM (Qu et al., 2022), and FedSWA (Caldarola et al., 2022)—and two recent fairness-without-disparity (FWD) approaches: KD-FedAvg (Chai et al., 2022) (adopted for FL settings) and FedSRCVaR (Papadaki et al., 2024).
5.1. Evaluation Datasets
Dataset Sensing Modality # of Participants ($N$) Application Task Sensitive Attributes
WIDAR (Zhang et al., 2021) WiFi (CSI Velocity Profiles) $N=12$ Gesture Recognition Sex, Sensor Orientation
HugaDB (Chereshnev and Kertész-Farkas, 2018) Inertial (Accel + Gyro) $N=18$ Human Activity Recognition Sex
Stress Sensing (Xiao et al., 2024) Physiological (EDA + Accel) $N=48$ Stress Detection Sex, Watch Hand
PERCEPT-R (Benway et al., 2022) Audio (Speech) $N=76$ Clinical Speech Assessment Sex
Table 1. Summary of datasets used for evaluation. The diversity in sensing modalities and tasks demonstrates the generalizability of CurvFed across the human-sensing spectrum.
We evaluate CurvFed on four heterogeneous human-sensing datasets, summarized in Table 1. These datasets were selected to test generalizability across diverse sensing modalities (WiFi, inertial, physiological, and audio) and distinct sensing tasks (HAR, stress detection, and speech therapy). Crucially, each dataset contains different sensitive attributes as biological sex, sensor placement, that challenge the fairness of standard FL algorithms (Yfantidou et al., 2023a; Djebrouni et al., 2024). In all our evaluation, the sensitive attributes were used solely for post-hoc fairness assessment.
WIDAR (WiFi-based HAR): This dataset leverages Channel State Information (CSI) derived from WiFi signals (Ratnam et al., 2024) to capture body-coordinate velocity profiles during human activities (Zhang et al., 2021). Data were collected from $16$ participants performing a set of gestures under varying conditions. To evaluate the robustness of CurvFed, we consider biological sex and receiver orientation as sources of natural variability across participants.
HugaDB (Inertial HAR): HugaDB (Chereshnev and Kertész-Farkas, 2018) focuses on wearable activity recognition and comprises accelerometer and gyroscope data collected from body-worn sensors on $18$ participants. The dataset provides a testbed for assessing whether CurvFed can achieve equitable performance across individuals with differing physiological characteristics, without relying on demographic information during training. For fairness evaluation purposes, biological sex is designated as the sensitive attribute.
Stress Sensing (Physiological): To evaluate multi-modal robustness, we utilize the Stress Sensing dataset (Xiao et al., 2024) ($48$ participants). It combines electrodermal activity (EDA) and accelerometer readings from an Empatica E4 wristband (McCarthy et al., 2016) to detect stress events. Sensitive attributes include biological sex and the hand (left vs. right) on which the device was worn.
PERCEPT-R (Clinical Audio): We extend our evaluation to the digital health domain using a publicly available subset of the PERCEPT corpus (v2.2.2) (Benway et al., 2022). The dataset consists of single-word speech recordings from children with speech sound disorders involving the rhotic / r / sound. In collaboration with clinical experts, we curated a cohort of $76$ participants and used expert-provided rhoticity annotations (0 = derhotic, 1 = fully rhotic). Sex is treated as the sensitive attribute to assess the equity of the clinical classification model across demographic groups. Further details on the datasets and FL client distributions for FL evaluations can be found in Appendix A.1.
5.2. Model Architectures and Hyper-parameters
We employ task-specific benchmark model architectures tailored to the computational constraints of edge devices and the specific modalities of each dataset.
HAR and Stress Sensing (MLPs): For the WIDAR, HugaDB, and Stress Sensing datasets, we adapt established benchmark architectures (Yang et al., 2023; Sharma et al., 2022; Pirttikangas et al., 2006). These are implemented as Multi-Layer Perceptrons (MLPs) with varying depth (2–3 layers) and regularization techniques (Dropout, Batch/Instance Normalization) optimized for their respective feature sets.
Clinical Audio (CNN): For the PERCEPT-R dataset, we utilized a Convolutional Neural Network (CNN) model (Benway et al., 2022) to process the temporal speech sequences. The network comprises two 1D convolutional layers (kernel size 5) that extract temporal features from the 5-channel input. The output is flattened and passed through a three-layer MLP classifier employing Hardswish activations to predict rhoticity.
Hyperparameters: We maintain consistent training budgets i.e. total number of FL rounds (R) and Local client training epochs (E) across all methods to ensure fair comparison. Table 2 details the specific hyperparameter settings for all four datasets. For the standard FedAvg, KD-Fedavg and FedSAM baselines, we utilize the "Base Learning Rate” and constant averaging. For CurvFed and FedSWA, we introduce the SWA components (SWA LR, $\alpha$, Cycle). Notably, the PERCEPT-R dataset requires a distinct configuration (lower $\alpha$, longer training) due to the higher complexity of the audio modality. Further details on grid search are provided in Appendix A.1.4.
Hyperparameter WIDAR Stress Sensing HugaDB PERCEPT-R
Training Budget
Local Epochs ($E$) 3 3 3 5
Total Rounds ($R$) 80 80 80 100
Optimization (CurvFed / Baseline)
Base Learning Rate ($\eta$) 0.01 / 0.1 0.01 / 0.1 0.001 / 0.001 0.001 / 0.001
SWA Learning Rate ($\eta_{swa}$) 0.1 0.001 0.001 0.001
Curvature Constraints (CurvFed)
Curvature Penalty ($\alpha$) 0.92 0.92 0.92 0.8
Cycle Length ($c$) 5 5 3 5
SWA Start Round 16 16 16 20
Stability Constant ($\epsilon$) (Eq 12) 0.005 0.005 0.005 0.005
Table 2. Hyperparameter configuration for main experiments. CurvFed introduces specific parameters (SWA, $\alpha$) to manage curvature. All methods share the same epoch and round budget to ensure fair comparison.